The joint probability density function of 𝑋 and π‘Œ is given by 𝑓(π‘₯, 𝑦) = { π‘₯ + 𝑦 3 0 ≀ π‘₯ ≀ 1, 0 ≀ 𝑦 ≀ 2 0 π‘’π‘™π‘ π‘’π‘€β„Žπ‘’π‘Ÿπ‘’ Find the two lines of regression.


1
Expert's answer
2021-12-15T05:14:04-0500

lines of regression:

"y=a_1x+b_1"

"a_1=\sigma_{XY}/\sigma^2_X,b_1=\mu_Y-a_1\mu_X"


"x=a_2y+b_2"

"a_2=\sigma_{XY}/\sigma^2_Y,b_2=\mu_X-a_2\mu_Y"


covariance:

"\sigma_{XY}=\iint (x-\mu_X)(y-\mu_Y)f(x,y)dydx"


"f_X(x)=\int^2_0 f(x,y)dy=\frac{1}{3}\int^2_0 (x+y)dy=\frac{1}{3}(2x+1)"


"f_Y(y)=\int^1_0 f(x,y)dx=\frac{1}{3}\int^1_0 (x+y)dx=\frac{1}{3}(1/2+y)"


"\mu_X=\int^1_0 xf_X(x)dx=\frac{1}{3}\int^1_0 x(2x+1)dx=\frac{1}{3}(2/3+1/2)=7/18"


"\mu_Y=\int^2_0 yf_Y(y)dy=\frac{1}{3}\int^2_0 y(1/2+y)dy=\frac{1}{3}(1+8/3)=11/9"


"\sigma_X=\sqrt{E(X^2)-\mu_X^2}"


"E(X^2)=\int^1_0 x^2f_X(x)dx=\frac{1}{3}\int^1_0 x^2(2x+1)dx=\frac{1}{3}(1/2+1/3)=5/18"


"\sigma_X=\sqrt{5/18-(7/18)^2}=\sqrt{41}/18=0.356"


"\sigma_Y=\sqrt{E(Y^2)-\mu_Y^2}"


"E(Y^2)=\int^2_0 x^2f_Y(y)dy=\frac{1}{3}\int^2_0 y^2(1/2+y)dy=\frac{1}{3}(4/3+4)=16/9"


"\sigma_Y=\sqrt{16/9-(11/9)^2}=\sqrt{23}/9=0.533"


"\sigma_{XY}=\frac{1}{3}\intop^1_0 \intop^2_0 (x-7/18)(y-11/9)(x+y)dydx="


"=\frac{1}{3}\intop^1_0(x-7/18)(xy^2/2-11xy/9+y^3/3-11y^2/18)|^2_0dx="


"=\frac{1}{3}\intop^1_0(x-7/18)(2x-22x/9+8/3-44/18)dx="


"=\frac{1}{3}\intop^1_0(x-0.39)(-0.44x+0.22)dx=\frac{1}{3}(-0.15+0.11+0.09-0.09)=-0.013"


"a_1=-0.013\/0.356^2=-0.1"

"b_1=11/9+0.1\cdot7/18=1.26"

"y=-0.1x+1.26"


"a_2=-0.013\/0.533^2=-0.05"

"b_1=7/18+0.05\cdot11/9=0.45"

"x=-0.05y+0.45"